OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -11x^4y^-2/55x^-5y^-1

OpenStudy (pfenn1):

\[\frac{-11x^4y^{-2}}{55x^{-5}y^{-1}}\]Remember \[x^{-a}=\frac{1}{x^a}\]

OpenStudy (pfenn1):

\[\frac {x^4}{x^{-5}}=x^4x^5=x^9\]\[-\frac{11}{55}=-\frac15\]\[\frac{y^{-2}}{y^{-1}}=\frac{y}{y^2}=\frac 1y\]So\[\frac{-11x^4y^{-2}}{55x^{-5}y^{-1}}=-\frac{x^9}{5y}\]